A box of mass 20 kilograms is kept on a smooth horizontal surface. Another object of mass 16 kg is kept on the box. The two objects are connected by light in extensible string and the string goes over smooth pulley . If a force of 6 newton is applied on the box , and the acceleration of the box.




1
Expert's answer
2022-03-21T17:57:28-0400

"a=\frac F{m_1}=0.3~\frac{m}{s^2},"

"T_1=m_1(g-a)=190~N,"

"T_2=m_2(g+a)=162~N."